Ammonia is contained in a piston-cylinder assembly. The ammonia undergoes a reversible, isobaric process from an initial pressure of 30 lbf/in2 and volume of 4 ft3/lb to a final temperature of 70° F. The mass of ammonia is 0.35 lb. Determine the amount of work done on the ammonia in Btu, the amount of heat transfer to the ammonia in Btu, and the temperature at which the heat transfer occurs in °F. Assume that the heat transfer occurs at a constant temperature. Neglect changes in kinetic energy and potential energy.

HInt: please use ammonia table for properties nit ideal gas equation. Here ammonia cannot be treated as ideal gas. If you answer wrong i will rate you negative
Answer #1


☺ Given that initial pressure P = 30 lbfline = 30 psi f ļ ļ feon initial volume da ufellb. ² final temperature Tz = 70°F Theh = 41.84 +0.02 4319 (569.20) 2 287.67 Blu/lb. his hfitx, Chfgi) hie Si= 0.0951 + 0.4319 (1.3349-0.0951) P=30 e Psi Si= 0.630=) The amount of heat transfer to ammonia m Chath) gor (0.35)( 650.19 - 287.67) PD 126.86 BLU mean Temperature at which heat